OpenStudy (anonymous):

A bike shop sells rolling strollers for runners with babies, with a total of 28 in the store. Some of the strollers have 3 wheels and some have 4. If there are 100 wheels, how many 3-wheel strollers are there?

OpenStudy (anonymous):

Well, the equation would be 3x + 4y = 100.

OpenStudy (texaschic101):

t = 3 wheeled strollers f = 4 wheeled strollers t + f = 28 ---- f = 28 - t 3t + 4f = 100 3t + 4(28 - t) = 100 3t + 112 - 4t = 100 3t - 4t = 100 - 112 -t = -12 t = 12 There is 12 three wheeled strollers in the store.